# Causal Discovery from Conditionally Stationary Time Series

### v3

@@ 1 Introduction
Deep learning has achieved profound success in vision and language modelling (Brown et al., 2020; Nichol et al., 2021). Still, it remains a grand challenge for deep neural networks to perform causal discovery (Yi et al., 2020; Girdhar & Ramanan, 2020; Sauer & Geiger, 2021), which is critical for interpretability, generalization, and robustness (Lake et al., 2017; Schölkopf et al., 2021). Theoretically, structure identifiability is key to ensure a unique correspondence between observations and the underlying causal structures (Peters et al., 2017). Practically, better algorithms are needed to accurately extract the causal structure from data.
In time series analysis, causal discovery identifies the underlying temporal causal structure of the observed sequences. Historical approaches rely on a restrictive assumption: stationarity (Granger, 1969; Peters et al., 2017; Tank et al., 2021; Löwe et al., 2022), while real-world data is often nonstationary with potential hidden confounders. To address this issue, three major strategies have been recently proposed: (1) modelling nonstationary noise (Huang et al., 2020; Gong et al., 2022); (2) introducing time-dependent effects with a fixed causal structure (Huang et al., 2015, 2019); and (3) using discrete latent variables to capture structural changes over time (Saggioro et al., 2020). Despite these advances, causal discovery in nonstationary time series under realistic assumptions remains an open challenge.
@@ 2.1 Causal discovery in stationary time series
where $\textbf{PA}^{(j)}(t-1)$ denotes the parents of $\bm{x}^{(j)}_{t}$ and $\bm{\epsilon}_{t}^{(j)}\sim p_{\bm{\varepsilon}}$ represents independent noise. In this setting, $\textbf{PA}^{(j)}(t-1)$ is constant in time and aligns with the summary graph $\mathcal{G}^{S}$.
@@ 3.3 Implementation
with a factorised prior on the edge labels $p_{\psi_{w}}(\mathcal{W})=\prod_{k=1}^{K}\prod_{ij}p_{\psi_{w}}(w_{ijk})$. We can guide $\psi_{w}$ though domain knowledge (e.g. sparsity). Given $\mathcal{W}\sim p_{\psi_{w}}(\mathcal{W})$ and $\bm{s}_{1:T}$, a sequence $\bm{x}_{1:T}$ is generated as
@@
$\log p(\bm{x}_{1:T})\geq-KL\Big{(}q_{\phi_{w}}(\mathcal{W}|\bm{x}_{1:T})\Big{|}\Big{|}p_{\psi_{w}}(\mathcal{W})\Big{)}\\ -\sum_{t=1}^{T}KL\Big{(}q_{\phi_{s}}(\bm{s}_{t}|\bm{x}_{1:T})\Big{|}\Big{|}p_{\psi_{s}}(\bm{s}_{t}|\bm{x}_{t:t-L_{x}},\bm{s}_{t-1:t-L_{s}})\Big{)}\\ +\mathbb{E}_{q_{\phi}(\mathcal{W},\bm{s}_{1:T}|\bm{x}_{1:T})}\Big{[}\log p_{\psi_{x}}(\bm{x}_{1:T}|\bm{s}_{1:T},\mathcal{W})\Big{]},$ (15)
where reparametrisation tricks (Jang et al., 2016) ensure backpropagation for discrete samples $\mathcal{W},\bm{s}_{1:T}\sim q_{\phi}(\mathcal{W},\bm{s}_{1:T}|\bm{x}_{1:T})$. To reduce variance, we use straight-through Gumbel-softmax and fix gradients to pass through unperturbed marginals (Ahmed et al., 2023). During training, Eq. (10) becomes:
@@ 4 Theoretical Considerations
As mentioned, SDCI considers samples $\bm{x}_{1:T}\sim\mathcal{D}$, each generated under different structures. The presented results remain valid, as identifiability holds even when considering the distribution of a single sample. However, the consistency of SDCI, along with other approaches within the same family, such as ACD (Löwe et al., 2022), remains an open challenge. Recent works Geffner et al. (2022); Gong et al. (2022) have shown that the validity of variational objectives for causal discovery relies on strong assumptions; including universal approximation properties of $q$, absence of model misspecification, and the infinite data limit. These works assume fixed structures across samples, where SDCI can be adapted by setting $q_{\phi}(\mathcal{W}|\bm{x}_{1:T})=q_{\phi}(\mathcal{W})$.
@@ 5 Related work
Our work focuses on modelling nonstationary time series with discrete latent variables that condition the underlying structure. It is related to Amortized Causal Discovery (ACD; Löwe et al. (2022)), which assumes stationarity and amortizes summary graph discovery across samples with different graphs but shared dynamics. We extend ACD by allowing the causal structure to vary according to latent state variables, thus enabling conditional stationarity. When observed auxiliary variables are present, Monti et al. (2020) uses time contrastive learning (Hyvarinen & Morioka, 2016) and nonlinear ICA for causal discovery, while Yao et al. (2022) establishes identifiability conditions in the presence of piece-wise stationary disturbances. Regime-dependent dynamics assume no access to states Saggioro et al. (2020), where Balsells-Rodas et al. (2024) provides identifiability results through Markov Switching Models. Varambally et al. (2024) also proves identifiability when the discrete latents represent global states. For states determined case, Song et al. (2024) establishes identifiability via mechanism sparsity and knowledge of the number of regimes. Other works focus on fixed causal structures, modelling nonstationarity through noise distributions (Huang et al., 2020; Gong et al., 2022) or time-varying effects (Huang et al., 2015; Zhang et al., 2017; Ghassami et al., 2018; Huang et al., 2019); such as Huang et al. (2020) with distribution shifts; Gong et al. (2022) with history-dependent noise; or Huang et al. (2015) with time-varying functional causal models.
@@ 6.1 Springs and Magnets
We evaluate SCDI on spring data adapted from Kipf et al. (2018); Löwe et al. (2022), consisting of particles in a box, connected by springs with directed impact - meaning that e.g. particle $i$ could affect particle $j$ with a force through a connecting spring, but leaving particle $i$ unaffected by this spring force. We further modify the dataset to introduce repulsive magnetic interactions, resulting in 3 edge-types: none, spring, and magnetic. We assess SDCI across determined and recurrent states. In the determined case, the state is modeled as a function of particles’ positions, while the recurrent case involves state transitions when particles collide with walls. Details on data generation, model hyper-parameters, and additional results (including observed case) are provided in Appendices C.1, B.4, and E.1 respectively.
@@
To empirically verify the consistency of SDCI in estimating state-dependent structures, we begin with a fixed-graph setting, where the graph remains constant across samples, and use an amortised posterior $q(\mathcal{W}|\bm{x}_{1:T}):=q(\mathcal{W})$ for inference. Fig. 4 demonstrates that SDCI achieves perfect structure estimation on both state settings. However, accurately estimating the states in the recurrent case is more challenging, likely due to the exponential computational cost of exact inference. Our formulation considers graphs that vary across samples, and we present results again in both state cases in Fig. 5. In this case, recovering state-dependent structures is significantly more challenging. While SDCI successfully captures the state dynamics in the determined case, the increased complexity from variable graph structures makes recovering state-dependent structures in the recurrent case notably difficult, even with only 3 variables.
We compare SDCI with causal discovery baselines in extracting graphs from conditionally stationary time series. Specifically, we measure the $F_{1}$ score for summary graph identification in determined states data with fixed and variable graph structures, benchmarking against ACD (Löwe et al., 2022), Rhino (Gong et al., 2022), and Neural Granger Causality (N-GC; (Tank et al., 2021)). For variable graph datasets, Rhino and N-GC require re-training for each sample, adding computational overhead. Results in Fig. 6 show a clear advantage of SDCI in terms of summary graph estimation. We observe that approaches assuming causal stationarity struggle to aggregate the full structure into a coherent summary graph, limiting their performance. However, N-GC particularly achieves strong results for variable graphs, surpassing ACD despite its simplicity. Overall, SDCI decomposes the nonstationary dynamics into conditional stationary components while accurately capturing state transition dynamics.
@@ 7 Conclusions
We have developed SDCI for causal discovery for conditionally stationary time series. The key innovation is the new concept of conditional summary graphs that are identifiable under a broad class of underlying state dependencies Evaluations on semi-synthetic data show SDCI’s superior performance in extracting the underlying causal graph and its potential for biology applications. Importantly, the improvement of SDCI over VRNN and NRI on modeling NBA player movements demonstrate the promise of causality-driven methods for forecasting and data interpretability.
@@ B.3 Encoder Architecture
Below we provide details our encoder arquitecture $q_{\phi}(\mathcal{W},\bm{s}_{1:T}|\bm{x}_{1:T})=q_{\phi_{w}}(\mathcal{W}|\bm{x}_{1:T})q_{\phi_{s}}(\bm{s}_{1:T}|\bm{x}_{1:T})=$.
@@
Our arquitecture extends directly from ACD (Löwe et al., 2022), and we incorporate additional edge-edge layers proposed in Chen et al. (2021) only for SDCI under the recurrent state case. The logits $\bm{\phi}_{ij}$ for the distribution $q_{\phi_{w}}(\mathcal{W}|\bm{x}_{1:T})$ are obtained as follows. First, the model computes a latent embedding $\bm{z}^{(i)}_{1}$ for each node $i$ using the whole sequence:
@@ B.4 Training specifications
Our method has been implemented with Pytorch (Paszke et al., 2019), and the experiments have been performed on NVIDIA RTX 2080 Ti GPUs.
@@
For GRN and springs, we tune the hyperparameters of Rhino based on the validation RMSE error. Our experimental setup follows a similar configuration to (Gong et al., 2022), with the key difference that we use a single MLP layer with 15 hidden units for the function $f_{i}$. We train the model using a batch size of 128 and an initial learning rate of 0.001, optimized with Adam. The node embedding dimensions are set to 32 and 16 for each respective dataset. Additionally, we use lag = 1, auglag = 60, and apply a sparsity penalty $\lambda_{s}$ of 19 and 25 for the two datasets, respectively.
@@ C.3 NBA Data
The NBA dataset11footnotetext: Data extracted from the following code repository https://github.com/linouk23/NBA-Player-Movements – last accessed 2022-09-28. (Linou, 2016) consists of recordings from 632 NBA games played during Winter 2015-2016. Each game is composed by approximately 400 to 600 events, which represents sequences of plays. In each trajectory, we find information about the ball location and 5 players of the 2 different teams (10 in total). The coordinates of the ball and players are represented in 3D and the length of each sequence can vary from 100 to 600. In our experiments, we consider a sequence up to $200$ time-steps ($T=100$ for reconstruction and the rest 100 steps for prediction), which gives us a total training dataset of 150K samples and a test set of 6380 samples. Data inspection shows that the court size is $100\times 50$, and we use this information to standardise the data. For experiments with SDCI with observed states, we design some ground-truth states which depend on different locations of the court. We set $K=4$ and our choice is shown in figure 12.

### v4

@@ 1 Introduction
Deep learning has achieved profound success in vision and language modelling (Brown et al., 2020; Nichol et al., 2022). Still, it remains a grand challenge for deep neural networks to perform causal discovery (Yi et al., 2020; Girdhar & Ramanan, 2020; Sauer & Geiger, 2021), which is critical for interpretability, generalization, and robustness (Lake et al., 2017; Schölkopf et al., 2021). Theoretically, structure identifiability is key to ensure a unique correspondence between observations and the underlying causal structures (Peters et al., 2017). Practically, better algorithms are needed to accurately extract the causal structure from data.
In time series analysis, causal discovery identifies the underlying temporal causal structure of the observed sequences. Historical approaches rely on a restrictive assumption: stationarity (Granger, 1969; Peters et al., 2017; Tank et al., 2021; Löwe et al., 2022), while real-world data is often nonstationary with potential hidden confounders. To address this issue, three major strategies have been recently proposed: (1) modelling nonstationary noise (Huang et al., 2020; Gong et al., 2023); (2) introducing time-dependent effects with a fixed causal structure (Huang et al., 2015, 2019); and (3) using discrete latent variables to capture structural changes over time (Saggioro et al., 2020). Despite these advances, causal discovery in nonstationary time series under realistic assumptions remains an open challenge.
@@ 2.1 Causal Discovery in Stationary Time Series
where $\textbf{PA}^{(j)}(t-1)$ denotes the parents [^1] The notation $(t-1)$ indicates that the parents of variable $j$ are considered at the previous time step. of $\bm{x}^{(j)}_{t}$ and $\bm{\epsilon}_{t}^{(j)}\sim p_{\bm{\varepsilon}}$ represents independent noise. In this setting, $\textbf{PA}^{(j)}(t-1)$ is constant in time and aligns with the summary graph $\mathcal{G}^{S}$.
@@ 3.3 Implementation
with a factorised prior on the edge labels $p_{\psi_{w}}(\mathcal{W})=\prod_{k=1}^{K}\prod_{ij}p_{\psi_{w}}(w_{ijk})$. We can guide $\psi_{w}$ through domain knowledge (e.g. sparsity). Given $\mathcal{W}\sim p_{\psi_{w}}(\mathcal{W})$ and $\bm{s}_{1:T}$, a sequence $\bm{x}_{1:T}$ is generated as
@@
$\log p_{\psi}(\bm{x}_{1:T})\geq-KL\Big{(}q_{\phi_{w}}(\mathcal{W}|\bm{x}_{1:T})\Big{|}\Big{|}p_{\psi_{w}}(\mathcal{W})\Big{)}\\ -\sum_{t=1}^{T}KL\Big{(}q_{\phi_{s}}(\bm{s}_{t}|\bm{x}_{1:T})\Big{|}\Big{|}p_{\psi_{s}}(\bm{s}_{t}|\bm{x}_{t:t-L_{x}},\bm{s}_{t-1:t-L_{s}})\Big{)}\\ +\mathbb{E}_{q_{\phi}(\mathcal{W},\bm{s}_{1:T}|\bm{x}_{1:T})}\Big{[}\log p_{\psi_{x}}(\bm{x}_{1:T}|\bm{s}_{1:T},\mathcal{W})\Big{]},$ (15)
where reparametrisation tricks (Jang et al., 2017) ensure backpropagation for discrete samples $\mathcal{W},\bm{s}_{1:T}\sim q_{\phi}(\mathcal{W},\bm{s}_{1:T}|\bm{x}_{1:T})$. To reduce variance, we use straight-through Gumbel-softmax and fix gradients to pass through unperturbed marginals (Ahmed et al., 2023). During training, Eq. (10) becomes:
@@ 4 Theoretical Considerations
While several prior works (Hälvä & Hyvarinen, 2020; Song et al., 2024; Balsells-Rodas et al., 2024) introduce nonstationarity via regime switching, our approach offers three key advantages:
  • Unknown number of state values $K$: Most prior work, including HMM-ICA (Hälvä & Hyvarinen, 2020) and CtrlNS (Song et al., 2024), assumes that the number of latent regimes is known. Our approach leverages Yakowitz & Spragins (1968), which enables identifiability without knowing the true number of regimes. This allows identification of $K$ by model selection (McLachlan & Peel, 2000) (assuming convergence to the MLE), which is not theoretically supported when $K$ requires to be known.
  • No assumption on state dependency: Prior works typically assume either the “states independent” case (Hälvä & Hyvarinen, 2020; Balsells-Rodas et al., 2024; Rahmani & Frossard, 2025); or the “states determined” case under strong structural constraints (Song et al., 2024). Our identifiability results make no assumptions about state dependencies, allowing feedback from observations.
  • Regime-dependent identifiability: Our theoretical results (Thms. A.6-A.7; Cor. A.8) build from general Markov Switching Models to our specific implementation, SDCI. Theorem A.6 introduces a novel proof strategy that extends from Yakowitz & Spragins (1968), providing a general theoretical foundation for regime-dependent causal discovery (Saggioro et al., 2020).
As mentioned, SDCI considers samples $\bm{x}_{1:T}\sim\mathcal{D}$, each generated under different structures. The presented results remain valid, as identifiability holds even when considering the distribution of a single sample. However, the consistency of SDCI, along with other approaches within the same family, such as ACD (Löwe et al., 2022), remains an open challenge. Recent works (Gong et al., 2023; Geffner et al., 2024) have shown that the validity of variational objectives for causal discovery relies on strong assumptions; including universal approximation properties of $q$, absence of model misspecification, and the infinite data limit. These works assume fixed structures across samples, where SDCI can be adapted by setting $q_{\phi_{w}}(\mathcal{W}|\bm{x}_{1:T})=q_{\phi_{w}}(\mathcal{W})$.
@@ 5 Related Work
Our work focuses on modelling nonstationary time series with discrete latent variables that condition the underlying structure. It is related to Amortized Causal Discovery (ACD; Löwe et al. (2022)), which assumes stationarity and amortizes summary graph discovery across samples with different graphs but shared dynamics. We extend ACD by allowing the causal structure to vary according to latent state variables, thus enabling conditional stationarity. When observed auxiliary variables are present, Monti et al. (2020) uses time contrastive learning (Hyvarinen & Morioka, 2016) and nonlinear ICA for causal discovery; while Yao et al. (2022) establishes identifiability conditions in the presence of piece-wise stationary disturbances. Regime-dependent dynamics assume no access to states Saggioro et al. (2020), where Balsells-Rodas et al. (2024) proves identifiability through Markov Switching Models; and Rahmani & Frossard (2025) via a score-based approach. Varambally et al. (2024) also proves identifiability when the discrete latents represent global states. Our “states determined” model is similar to Song et al. (2024), where identifiability is established by assuming mechanism sparsity and knowledge of the number of regimes. Other works focus on fixed causal structures, modelling nonstationarity through noise distributions (Huang et al., 2020; Gong et al., 2023) or time-varying effects (Huang et al., 2015; Zhang et al., 2017; Ghassami et al., 2018; Huang et al., 2019); such as Huang et al. (2020) with distribution shifts; Gong et al. (2023) with history-dependent noise; or Huang et al. (2015) with time-varying functional causal models.
@@ 6.1 Springs and Magnets
We evaluate SCDI on spring data adapted from Kipf et al. (2018); Löwe et al. (2022), consisting of particles in a box, connected by springs with directed impact – meaning that e.g. particle $i$ could affect particle $j$ with a force through a connecting spring, but leaving particle $i$ unaffected by this spring force. We further modify the dataset to introduce repulsive magnetic interactions, resulting in 3 edge-types: none, spring, and magnetic. We assess SDCI across determined and recurrent states. In the determined case, the state is modeled as a function of particles’ positions, while the recurrent case involves state transitions when particles collide with walls. Details on data generation, model hyper-parameters, and additional results (including observed case) are provided in Appendices C.1, B.4, and E.1 respectively.
@@
To empirically verify the consistency of SDCI in estimating state-dependent structures, we begin with a fixed-graph setting, where the graph remains constant across samples, and use an amortised posterior $q_{\phi_{w}}(\mathcal{W}|\bm{x}_{1:T})=q_{\phi_{w}}(\mathcal{W})$ for inference. Fig. 4 demonstrates that SDCI achieves perfect structure estimation on both state settings. However, accurately estimating the states in the recurrent case is more challenging, likely due to the exponential computational cost of exact inference. Our formulation considers graphs that vary across samples, and we present results again in both state cases in Fig. 5. In this case, recovering state-dependent structures is significantly more challenging. While SDCI successfully captures the state dynamics in the determined case, the increased complexity from variable graph structures makes recovering state-dependent structures in the recurrent case notably difficult, even with only 3 variables.
We compare SDCI with causal discovery baselines in extracting graphs from conditionally stationary time series. Specifically, we measure the $F_{1}$ score for summary graph identification in determined states data with fixed and variable graph structures, benchmarking against ACD (Löwe et al., 2022), Rhino (Gong et al., 2023), and Neural Granger Causality (N-GC; (Tank et al., 2021)). For variable graph datasets, Rhino and N-GC require re-training for each sample, adding computational overhead. Results in Fig. 6 show a clear advantage of SDCI in terms of summary graph estimation. We observe that approaches assuming causal stationarity struggle to aggregate the full structure into a coherent summary graph, limiting their performance. However, N-GC particularly achieves strong results for variable graphs, surpassing ACD despite its simplicity. Overall, SDCI decomposes the nonstationary dynamics into conditional stationary components while accurately capturing state transition dynamics.
Selecting the number of states.
@@
Figure 7: Test reconstruction MSE vs number of states $K$ for data with ground truth $K=2$ (red) and $K=4$ (blue).
Determining the number of states $K$, or the number of regimes in regime-switching dynamics, is a fundamental challenge in real-world data. As discussed in Section 4, SDCI is also identifiable in terms of $K$. Assuming the consistency of SDCI, $K$ can be theoretically recovered by maximising the data log-likelihood, which we approach by model selection (McLachlan & Peel, 2000). However, in practice, overestimating $K$ may lead to overfitting or redundant regimes. To overcome this issue, standard selection heuristics such as the elbow method (Thorndike, 1953) can be applied. Figure 7 shows the reconstruction error for two settings: states recurrent with $N=5$ and $K=2$ (in red); and states determined with $N=10$ and $K=4$ (in blue). In both cases, the MSE plateaus when $K$ reaches its ground truth value, thus providing a principled heuristic to identify the true number of states.
@@ 7 Conclusions
We have developed SDCI for causal discovery in conditionally stationary time series. The key innovation is the new concept of conditional summary graphs that are identifiable under a broad class of underlying state dependencies Evaluations on semi-synthetic data show SDCI’s superior performance in extracting the underlying causal graph and its potential for biology applications. Importantly, the improvement of SDCI over VRNN and NRI on modeling NBA player movements demonstrate the promise of causality-driven methods for forecasting and data interpretability.
@@ A.3 Sufficiency Conditions for Identifiability
$\mathcal{X}_{e,e^{\prime}}:=\bigg{\{}\bm{x}_{1}\in\mathbb{R}^{d}:\exists\bm{x}_{2}\in\mathbb{R}^{d},\quad\frac{\partial f_{e}(\bm{x}_{1},\bm{x}_{2})}{\partial\bm{x}_{1}}=\frac{\partial f_{e}^{\prime}(\bm{x}_{1},\bm{x}_{2})}{\partial\bm{x}_{1}}\bigg{\}}.$ (33)
@@ A.4 Identifiability of Markov Switching Models
To clarify, Proposition 2 in Yakowitz & Spragins (1968) establishes multivariate Gaussian distributions are identifiable if and only if the means or covariances are distinct. Our assumptions (b1) and (b2) ensure distinct means for almost any $\bm{x}_{t-1}=\bm{\alpha}\in\mathbb{R}^{Nd}$.
@@ B.3 Encoder Architecture
Below we provide details our encoder architecture $q_{\phi}(\mathcal{W},\bm{s}_{1:T}|\bm{x}_{1:T})=q_{\phi_{w}}(\mathcal{W}|\bm{x}_{1:T})q_{\phi_{s}}(\bm{s}_{1:T}|\bm{x}_{1:T})$.
@@
Our architecture extends directly from ACD (Löwe et al., 2022), and we incorporate additional edge-edge layers proposed in Chen et al. (2021) only for SDCI under the recurrent state case. The logits $\bm{\phi}_{ij}$ for the distribution $q_{\phi_{w}}(\mathcal{W}|\bm{x}_{1:T})$ are obtained as follows. First, the model computes a latent embedding $\bm{z}^{(i)}_{1}$ for each node $i$ using the whole sequence:
@@ B.4 Training Specifications
Our method has been implemented with Pytorch (Paszke et al., 2019), and the experiments have been performed on NVIDIA RTX 2080 Ti GPUs. Code for our experiments is available at https://github.com/charlio23/SDCI.
@@
For GRN and springs, we tune the hyperparameters of Rhino based on the validation RMSE error. Our experimental setup follows a similar configuration to (Gong et al., 2023), with the key difference that we use a single MLP layer with 15 hidden units for the function $f_{i}$. We train the model using a batch size of 128 and an initial learning rate of 0.001, optimized with Adam. The node embedding dimensions are set to 32 and 16 for each respective dataset. Additionally, we use lag = 1, auglag = 60, and apply a sparsity penalty $\lambda_{s}$ of 19 and 25 for the two datasets, respectively.
@@ C.3 NBA Data
The NBA dataset22footnotetext: Data extracted from the following code repository https://github.com/linouk23/NBA-Player-Movements – last accessed 2022-09-28. (Linou, 2016) consists of recordings from 632 NBA games played during Winter 2015-2016. Each game is composed by approximately 400 to 600 events, which represents sequences of plays. In each trajectory, we find information about the ball location and 5 players of the 2 different teams (10 in total). The coordinates of the ball and players are represented in 3D and the length of each sequence can vary from 100 to 600. In our experiments, we consider a sequence up to $200$ time-steps ($T=100$ for reconstruction and the rest 100 steps for prediction), which gives us a total training dataset of 150K samples and a test set of 6380 samples. Data inspection shows that the court size is $100\times 50$, and we use this information to standardise the data. For experiments with SDCI with observed states, we design some ground-truth states which depend on different locations of the court. We set $K=4$ and our choice is shown in figure 14.